# Daily-Omni: Towards Audio-Visual Reasoning with Temporal Alignment across Modalities

### v1

@@ Abstract
Recent Multimodal Large Language Models (MLLMs) achieve promising performance on visual and audio benchmarks independently. However, the ability of these models to process cross-modal information synchronously remains largely unexplored. In this paper, we introduce: 1) Daily-Omni, an Audio-Visual Questioning and Answering benchmark comprising 684 videos of daily life scenarios from diverse sources, rich in both audio and visual information, and featuring 1197 multiple-choice QA pairs across 6 major tasks; 2) Daily-Omni QA Generation Pipeline, which includes automatic annotation, QA generation and QA optimization, significantly improves efficiency for human evaluation and scalability of the benchmark; 3) Daily-Omni-Agent, a training-free agent utilizing open-source Visual Language Model (VLM), Audio Language Model (ALM) and Automatic Speech Recognition (ASR) model to establish a baseline for this benchmark. The results show that current MLLMs still struggle significantly with tasks requiring audio-visual integration, but combining VLMs and ALMs with simple temporal alignment techniques can achieve substantially better performance. Codes and benchmark are available at https://github.com/Lliar-liar/Daily-Omni.
@@ 1 Introduction
Non-textual modalities, such as vision and hearing, often convey richer and more nuanced information than text alone in daily life scenarios, playing a crucial role in our understanding of and interaction with the physical world. Therefore, advancements in Multimodal Large Language Models (MLLMs) capable of comprehensively understanding the multi-modal information are a crucial foundation for achieving artificial intelligence capable of interacting with the physical world and observing the outcome of its operations.
@@
Recent MLLMs [xu_qwen25-omni_2025, sun_video-salmonn_2024, noauthor_introducing_2024, zhang_internlm-xcomposer25-omnilive_2024, fu_vita-15_2025, liu_ola_2025, cheng_videollama_2024, li_baichuan-omni-15_2025, xie_mini-omni2_2024, guo_m2-omni_2025, team_reka_2024, microsoft2025phi4minitechnicalreportcompact, lu2023unifiedio2scalingautoregressive, liu_nexus-o_2025, gemmateam2025gemma3technicalreport] have demonstrated groundbreaking capabilities spanning audio (ASR, sound classification, captioning) and visual domains (OCR, VQA, video grounding), with significant accuracy improvements over previous benchmarks. However, existing methods still face several limitations. Firstly, many MLLMs predominantly focus on visual abilities, often neglecting the importance of other modalities like audio. This oversight may stem from the fact that current visual datasets are more abundant, of higher quality, and cover a broader range of tasks compared to audio datasets. Existing audio datasets [Librispeech, wang2020covost2massivelymultilingual, poria2019meldmultimodalmultipartydataset, chen_vggsound_2020, Audioset, Vocalsound] tend to prioritize speech-related or music-related tasks and basic sound classification, often overlooking more complex yet crucial tasks such as reasoning over generic sounds. Consequently, many MLLMs incorporate only speech encoders as their primary auditory component or rely heavily on speech-related datasets for audio pre-training. This architectural limitation fundamentally restricts their ability to comprehend rich acoustic environments where non-speech sounds (e.g., environmental noises, mechanical failures, or emotional cues in non-verbal vocalizations) carry critical semantic information. Secondly, the current landscape lacks high-quality multimodal datasets that integrate temporally aligned auditory and visual information. Existing audio-visual datasets and benchmarks [yun_pano-avqa_2021, li_learning_2022, avqa, li_omnibench_2024, hong_worldsense_2025, gong_av-odyssey_2024, sungbin2025avhbenchcrossmodalhallucinationbenchmark, geng2025longvalevisionaudiolanguageeventbenchmarktimeaware] reveal three persistent limitations. First, several focus on specialized scenarios [yun_pano-avqa_2021, li_learning_2022] such as musical performances or panoramic environments, thereby introducing domain-specific biases. Second, many employ static image-audio pairs [li_omnibench_2024, gong_av-odyssey_2024] that disregard crucial temporal dynamics inherent in real-world video contexts. Third, current tasks are often too narrow, with many benchmarks focusing on specific applications such as captioning or open-ended responses [geng2025longvalevisionaudiolanguageeventbenchmarktimeaware]. The common absence of rigorous evaluation frameworks and standardized metrics makes it hard to reliably compare results. While WorldSense [hong_worldsense_2025] established a valuable audio-visual multi-choice QA benchmark for daily scenarios through its meticulous dataset curation, two critical limitations persist: (1) it lacks a systematic framework for scalable QA generation, relying instead on manual annotation processes that hinder dataset expansion; and (2) the benchmark serves primarily as an evaluation tool, offering limited methodological guidance on enhancing model capabilities via explicit training protocols or architectural modifications. This dual limitation constrains both the benchmark’s adaptability to emerging domains and its practical utility in driving model improvements.
Figure 1: Examples of Daily-Omni QAs. The audio and visual information required for answering the questions are provided in the figure. The correct answer for the given questions are highlighted.
In this paper, we introduce Daily-Omni, an Audio-Visual Questioning and Answering benchmark with 684 videos featuring daily life scenarios from various sources [Audioset, fu2024videommefirstevercomprehensiveevaluation, Farré2024FineVideo] with rich audio and visual information and 1197 multiple choice QA-pairs across 6 major tasks ranging from audio visual event aligning to complicated cross-modal reasoning. The videos are sampled from different datasets and segmented into 30-second or 60-second intervals to systematically evaluate model performance across different temporal contexts. We further introduce Daily-Omni QA Generation Pipeline which encompasses five automated modules: video annotation, annotation revision, audio-visual temporal alignment, QA generation, and QA optimization. This framework demonstrates remarkable scalability: a single annotator can complete quality filtering process for these 1197 high-quality QA pairs within 30 hours, achieving an approximate 30% acceptance rate from initially generated candidates. In addition, we propose Daily-Omni Agent, an audio-visual agent utilizing open-source Visual Language Model (VLM), Audio Language Model (ALM) and Automatic Speech Recognition (ASR) model without further finetuning to establish a baseline for this benchmark. We evaluated recent MLLMs and our agent on the Daily-Omni benchmark, where the Daily-Omni Agent achieved state-of-the-art performance among open-source methods. The experimental results show that current MLLMs still face significant challenges in tasks requiring deep audio-visual temporal integration. They also reveal that by combining existing visual and audio language models with simple temporal alignment techniques, as demonstrated by our Daily-Omni Agent, substantially improved performance can be achieved, underscoring a promising direction for enhancing multimodal reasoning.
@@ 2.1 Multimodal Large Language Models
Recent advancements in Multimodal Large Language Models (MLLMs) primarily fall into three categories: Audio Language Models (ALMs) incorporating audio capabilities [tang_salmonn_2024, gong2024listenthinkunderstand, chu_qwen-audio_2023, chu2024qwen2audiotechnicalreport, ghosh2025audioflamingo2audiolanguage, ghosh2024gamalargeaudiolanguagemodel], Visual Language Models (VLMs) adding visual understanding [liu2025nvilaefficientfrontiervisual, bai_qwen25-vl_2025, lu2024deepseekvlrealworldvisionlanguageunderstanding, wang2024emu3nexttokenpredictionneed], and Omni-modal Language Models (OLMs) combining both audio and visual modalities [noauthor_introducing_2024, li_baichuan-omni-15_2025, zhang_internlm-xcomposer25-omnilive_2024, team_reka_2024, cheng_videollama_2024, liu_ola_2025, microsoft2025phi4minitechnicalreportcompact, fu_vita-15_2025, guo_m2-omni_2025, xu_qwen25-omni_2025, liu_nexus-o_2025, gemmateam2025gemma3technicalreport]. These MLLMs often employ a modular architecture, utilizing separate encoders for audio and visual inputs. In the audio domain, encoders such as [radford2022robustspeechrecognitionlargescale, chen_beats_2022, li2024mertacousticmusicunderstanding] are commonly used to extract sound representations. Some models, such as [tang_salmonn_2024, liu_ola_2025, zhang_internlm-xcomposer25-omnilive_2024], even integrate multiple audio encoders specialized for different sound types (e.g., speech, music). Similarly, visual encoders [liu2021swintransformerhierarchicalvision, li2023blip2bootstrappinglanguageimagepretraining, dai2023instructblipgeneralpurposevisionlanguagemodels, liu2023visualinstructiontuning, dehghani2023patchnpacknavit] are used to process images and videos. However, this modular approach, while cost-efficient and effective for individual modalities, struggles to capture the crucial temporal correlations inherent in synchronized audio-visual streams. This limitation arises because audio and visual inputs are encoded independently. Although multi-modal positional embedding techniques like M-RoPE [bai2023qwenvlversatilevisionlanguagemodel] and TMRoPE [xu_qwen25-omni_2025] enhance cross-modal temporal understanding to some extent, effective methods for deeply integrating and temporally aligning multimodal data remain relatively scarce. Additionally, in several MLLMs [microsoft2025phi4minitechnicalreportcompact, fu_vita-15_2025, zhang_internlm-xcomposer25-omnilive_2024] , the audio encoder primarily serves to process user instructions–akin to how the text modality is used–rather than sensing the ambient environment, which is a common function of the visual modality.
@@ 2.2 Audio-Visual Understanding Datasets and Benchmarks
The development of uni-modal datasets and associated tasks for audio and vision has driven advancements in Audio Language Models (ALMs) and Visual Language Models (VLMs). Visual datasets and benchmarks [li2024llavanextinterleavetacklingmultiimagevideo, fu2024videommefirstevercomprehensiveevaluation, wu_star_2024, mangalam_egoschema_2023, liu2023visualinstructiontuning, wu_longvideobench_nodate, zhang2024videoinstructiontuningsynthetic, yue2024mmmumassivemultidisciplinemultimodal, yue2024mmmuprorobustmultidisciplinemultimodal, li2023seedbenchbenchmarkingmultimodalllms, fu2024ocrbenchv2improvedbenchmark, gao2017talltemporalactivitylocalization, hu2025videommmuevaluatingknowledgeacquisition] primarily focus on tasks for static images (such as OCR, grounding, segmentation, classification, and question-answering) and dynamic videos (such as captioning, temporal grounding, and understanding), while audio datasets and benchmarks [ghosh2025audioflamingo2audiolanguage, Audioset, chen_vggsound_2020, Librispeech, yang2024airbenchbenchmarkinglargeaudiolanguage] address speech-related tasks (including ASR, emotion recognition, and entity recognition) and non-speech tasks (such as sound classification and audio grounding). However, while attempts at creating audio-visual datasets date back to at least 2021 [yun_pano-avqa_2021], these early efforts often had significant limitations. For example, Music-AVQA [li2022learninganswerquestionsdynamic] focused specifically on music performance videos, while Pano-AVQA [yun_pano-avqa_2021] centered on panoramic videos. Others, such as AVQA [avqa] and OmniBench [li_omnibench_2024], were restricted to short, simple videos or static images. Furthermore, AV-Odyssey [gong_av-odyssey_2024] heavily emphasized specific audio tasks, such as recognizing timbre and loudness, rather than broader audio-visual understanding. While WorldSense [hong_worldsense_2025] provides a valuable model for real-world audio-visual question-answering, how to efficiently scale AVQA datasets and enhance OLM abilities is an area requiring further exploration.
@@ 3 Daily-Omni
This section details the Daily-Omni framework, which comprises three core components: multi-modal data curation, hierarchical video annotation, and QA synthesis and evaluation. Furthermore, we delineate the evaluation paradigm implemented through the Daily-Omni Agent, which serves as our benchmarking baseline for systematic assessment. Daily-Omni aims to establish a benchmarking framework designed to systematically evaluate MLLMs’ ability to perceive and reason in real-life audio-visual scenarios. As illustrated in Figure LABEL:Fig:qa_count, the Daily-Omni benchmark comprises videos from all 11 YouTube categories. The QAs are meticulously crafted to assess a spectrum of multi-modal capabilities, spanning from basic cross-modal perception to intricate reasoning. All questions are designed to require the integration of audio, visual, and textual information for a correct answer. In total, the Daily-Omni benchmark comprises 684 videos and 1197 QA pairs. Of these QAs, 550 correspond to 60-second videos and 647 correspond to 30-second videos. We compared the features of Daily-Omni with various audio-visual benchmarks in Table 1.
@@ 3.1 Data Curation
Daily-life scenarios offer abundant information across both audio and visual modalities. Within the visual modality, our data curation focuses on selecting videos characterized by significant temporal dynamics. Conversely, static scenes, exemplified by a vlogger addressing the camera with limited motion, offer minimal temporal information for questioning and are usually excluded. To prevent data bias, videos are selected from a wide range of categories and topics. Videos with harmful or discriminative content are excluded. Regarding the audio modality, recognizing that prior datasets have concentrated heavily on speech, our benchmark is designed to encompass a broader range of everyday sounds, such as music, speech, and other sound events. These acoustic signals should be present within the videos, appearing either concurrently or consecutively. Furthermore, to maintain focus on audio-visual processing and avoid challenges related to multilingual text comprehension, videos with languages other than English are also excluded.
Our video data originates from AudioSet [Audioset], Video-MME [fu2024videommefirstevercomprehensiveevaluation], and FineVideo [Farré2024FineVideo]. While AudioSet provides 10-second clips primarily for audio classification, we retrieved the original, full-length source videos. These were then processed into longer 30s or 60s segments which contains the original 10s segments to ensure they contain certain types of sound events. To ensure inclusion of spoken content, we employed Whisper-V3-Large [radford2022robustspeechrecognitionlargescale] to detect speech within these segments, adding those segments identified as containing speech to our set of video candidates. For Video-MME and FineVideo, our selection criteria prioritized videos containing substantial audio information alongside rich temporal visual dynamics. The inclusion of these datasets also increases diversity, as they feature more recent videos compared to Audio Set (primarily pre-2017 uploads), thus broadening the representation of genres and styles.

### v2

@@ Abstract
Recent Multimodal Large Language Models (MLLMs) achieve promising performance on visual and audio benchmarks independently. However, the ability of these models to process cross-modal information synchronously remains largely unexplored. We introduce Daily-Omni, a multiple-choice Audio-Visual QA benchmark featuring 684 real-world videos and 1,197 questions spanning 6 task families that explicitly require cross-modal temporal reasoning. To support scalable benchmark construction, we develop a semi-automatic pipeline for annotation, cross-modal consistency refinement, temporal alignment elicitation, and text-only leakage filtering, followed by human verification. We further provide a diagnostic evaluation suite and extensively evaluate 24 foundation models under 37 model–modality settings (Audio+Video / Audio-only / Video-only / Text-only). Finally, we include a training-free modular diagnostic baseline that composes off-the-shelf unimodal models to serve as a diagnostic baseline and to illustrate how explicit temporal alignment signals affect performance. Results indicate that many end-to-end MLLMs still struggle on alignment-critical questions, suggesting that robust cross-modal temporal alignment remains an important open challenge.
@@ 1 Introduction
In our daily lives, non-textual modalities like vision and audio convey much richer information than text alone. They are essential for us to understand and interact with the physical world. Therefore, advancements in Multimodal Large Language Models (MLLMs) capable of comprehensively understanding the multi-modal information are a crucial foundation for achieving artificial intelligence capable of interacting with the physical world and observing the outcome of its operations.
@@
Recent MLLMs [Qwen3-VL, Qwen3-Omni, xu_qwen25-omni_2025, tang_enhancing_2024, noauthor_introducing_2024, zhang_internlm-xcomposer25-omnilive_2024, fu_vita-15_2025, liu_ola_2025, cheng_videollama_2024, li_baichuan-omni-15_2025, xie_mini-omni2_2024, guo_m2-omni_2025, team_reka_2024, microsoft2025phi4minitechnicalreportcompact, lu2023unifiedio2scalingautoregressive, liu_nexus-o_2025, gemmateam2025gemma3technicalreport, gemini25] have demonstrated groundbreaking capabilities spanning audio (ASR, sound classification, captioning) and visual domains (OCR, VQA, video grounding), with significant accuracy improvements over previous benchmarks. However, existing methods still face several limitations. Firstly, many MLLMs predominantly focus on visual abilities, often neglecting the importance of other modalities like audio. This oversight may stem from the fact that current visual datasets are more abundant, of higher quality, and cover a broader range of tasks compared to audio datasets. Existing audio datasets [Librispeech, wang2020covost2massivelymultilingual, poria2019meldmultimodalmultipartydataset, chen_vggsound_2020, Audioset, Vocalsound] tend to prioritize speech-related or music-related tasks and basic sound classification, often overlooking more complex yet crucial tasks such as reasoning over generic sounds. Consequently, many MLLMs incorporate only speech encoders as their primary auditory component or rely heavily on speech-related datasets for audio pre-training. This architectural limitation fundamentally restricts their ability to comprehend rich acoustic environments where non-speech sounds (e.g., environmental noises, mechanical failures, or emotional cues in non-verbal vocalizations) carry critical semantic information. Secondly, the current landscape lacks high-quality multimodal datasets that integrate temporally aligned auditory and visual information. Existing audio-visual datasets and benchmarks [yun_pano-avqa_2021, li2022learninganswerquestionsdynamic, avqa, li_omnibench_2024, hong_worldsense_2025, gong_av-odyssey_2024, sungbin2025avhbenchcrossmodalhallucinationbenchmark, geng2025longvalevisionaudiolanguageeventbenchmarktimeaware, omnivideobench] suffer from three persistent limitations. First, several focus on specialized scenarios [yun_pano-avqa_2021, li2022learninganswerquestionsdynamic] such as musical performances or panoramic environments, thereby introducing domain-specific biases.
Figure 1: Illustrative examples from the Daily-Omni benchmark. These cases highlight the dataset’s focus on complex cross-modal reasoning. The top example requires temporally aligning specific speakers’ voices with their gestures, while the bottom example requires comparing acoustic environments across different video segments.
Second, many employ static image-audio pairs [li_omnibench_2024, gong_av-odyssey_2024] that disregard crucial temporal dynamics inherent in real-world video contexts. Third, current tasks are often too narrow, with many benchmarks focusing on specific applications such as captioning or open-ended responses [geng2025longvalevisionaudiolanguageeventbenchmarktimeaware]. The common absence of rigorous evaluation frameworks and standardized metrics makes it hard to reliably compare results. While WorldSense [hong_worldsense_2025] established a valuable audio-visual multi-choice QA benchmark for daily scenarios through its meticulous dataset curation, two critical limitations persist: (1) it lacks a systematic framework for scalable QA generation, relying instead on manual annotation processes that hinder dataset expansion; and (2) the benchmark serves primarily as an evaluation tool, offering limited methodological guidance on enhancing model capabilities via explicit training protocols or architectural modifications. This dual limitation constrains both the benchmark’s adaptability to emerging domains and its practical utility in driving model improvements.
In this paper, we introduce Daily-Omni, an Audio-Visual Questioning and Answering benchmark designed to evaluate temporally-aligned multimodal reasoning. The videos are sampled from diverse datasets [Audioset, fu2024videommefirstevercomprehensiveevaluation, Farre2024FineVideo] and segmented into 30-second or 60-second intervals to systematically assess model performance across different temporal contexts. Our core contributions are three-fold:
  • The Daily-Omni Benchmark: We construct a high-quality dataset comprising 684 real-life videos rich in both audio and visual dynamics, featuring 1197 multiple-choice QA pairs across 6 major tasks ranging from fundamental audio-visual event alignment to complicated cross-modal reasoning.
  • Scalable QA Generation Pipeline: We propose an automated framework encompassing video annotation, revision, audio-visual temporal alignment, QA generation, and optimization. This highly efficient pipeline significantly reduces human labor, requiring only 30 hours for a single annotator to complete the quality filtering process (achieving an approximate 30% acceptance rate), thereby ensuring the benchmark’s scalability.
  • Extensive Evaluation & Diagnostic Protocol. We benchmark 24 representative foundation models under 37 model–modality settings (AV / V-only / A-only / text-only) and provide diagnostic analyses including modality ablations and alignment-sensitivity. We additionally include a training-free modular reference baseline (Daily-Omni Agent) composed of off-the-shelf unimodal models, intended as a sanity-check reference.
Our comprehensive results reveal an important phenomenon: many current end-to-end MLLMs struggle with tasks requiring deep audio-visual temporal integration, to the extent that our simple, explicit-alignment baseline outperforms several recent open-source omni-modal models. This extensive evaluation demonstrates that robust temporal alignment mechanisms are still insufficient or missing in many unified architectures, underscoring both the necessity of our benchmark and a clear direction for future multimodal reasoning research.
@@ 2.1 Multimodal Large Language Models
Recent advancements in Multimodal Large Language Models (MLLMs) primarily fall into three categories: Audio Language Models (ALMs) incorporating audio capabilities [tang_salmonn_2024, gong2024listenthinkunderstand, chu_qwen-audio_2023, chu2024qwen2audiotechnicalreport, ghosh2025audioflamingo2audiolanguage, ghosh2024gamalargeaudiolanguagemodel, goel2025audioflamingo3advancing], Visual Language Models (VLMs) adding visual understanding [liu2025nvilaefficientfrontiervisual, bai_qwen25-vl_2025, lu2024deepseekvlrealworldvisionlanguageunderstanding, wang2024emu3nexttokenpredictionneed], and Omni-modal Language Models (OLMs) combining both audio and visual modalities [noauthor_introducing_2024, li_baichuan-omni-15_2025, zhang_internlm-xcomposer25-omnilive_2024, team_reka_2024, cheng_videollama_2024, liu_ola_2025, microsoft2025phi4minitechnicalreportcompact, fu_vita-15_2025, guo_m2-omni_2025, Qwen3-Omni, xu_qwen25-omni_2025, liu_nexus-o_2025, gemmateam2025gemma3technicalreport]. These MLLMs often employ a modular architecture, utilizing separate encoders for audio and visual inputs. In the audio domain, several works [chen2022beatsaudiopretrainingacoustic, li2024mertacousticmusicunderstanding, radford2022robustspeechrecognitionlargescale] proposed audio encoders to extract sound representations. Some models [tang_salmonn_2024, liu_ola_2025, zhang_internlm-xcomposer25-omnilive_2024] even integrate multiple audio encoders specialized for different sound types (e.g., speech, music). Similarly, visual encoders [li2023blip2bootstrappinglanguageimagepretraining, dai2023instructblipgeneralpurposevisionlanguagemodels, liu2023visualinstructiontuning, dehghani2023patchnpacknavit] are used to process images and videos. However, this modular approach struggles to capture the crucial temporal correlations inherent in synchronized audio-visual streams. This limitation arises because audio and visual inputs are encoded independently. Although multi-modal positional embedding techniques like TMRoPE [xu_qwen25-omni_2025] enhance cross-modal temporal understanding to some extent, effective methods for temporally aligning multimodal data remain relatively scarce. Additionally, in several MLLMs [microsoft2025phi4minitechnicalreportcompact, fu_vita-15_2025, zhang_internlm-xcomposer25-omnilive_2024] , the audio encoder primarily serves to process user instructions—akin to how the text modality is used—rather than perceiving the environment.
Table 1: Comparison of audio-visual benchmarks. We detail publication, modality (A: audio, V: video, I: image), size, and question type (MC: multiple choice, DF: defined word, $B$: bounding boxes). ’Automatic Scalability’ refers to automated expansion methods. ’Open-Domain’ and ’General Sound’ denote genre and sound diversity.
Benchmarks Pub Modality #QA Question Type Automatic Scalability Open-Domain General Sound
AVQA ACM MM’22 V+A 57,335 MC
Music-AVQA CVPR’22 V+A 45,867 DF
Pano-AVQA ICCV’21 V+A 51,700 DF & BB
OmniBench ARXIV’24 I+A 1,142 MC
AV-Odyssey ARXIV’24 I+A 4,555 MC
WorldSense ICLR’26 V+A 3,172 MC
Daily-Omni V+A 1197 MC
@@ 2.2 Audio-Visual Understanding Datasets and Benchmarks
The development of uni-modal datasets and associated tasks for audio and vision has driven advancements in Audio Language Models (ALMs) and Visual Language Models (VLMs). Visual datasets and benchmarks [li2024llavanextinterleavetacklingmultiimagevideo, fu2024videommefirstevercomprehensiveevaluation, wu_star_2024, mangalam_egoschema_2023, liu2023visualinstructiontuning, wu2024longvideobench, zhang2024videoinstructiontuningsynthetic, yue2024mmmumassivemultidisciplinemultimodal, yue2024mmmuprorobustmultidisciplinemultimodal, li2023seedbenchbenchmarkingmultimodalllms, fu2024ocrbenchv2improvedbenchmark, gao2017talltemporalactivitylocalization, hu2025videommmuevaluatingknowledgeacquisition, omnivideobench] primarily focus on tasks for static images (OCR, grounding, segmentation, classification, and question-answering) and dynamic videos (captioning, temporal grounding, and understanding), while audio datasets and benchmarks [ghosh2025audioflamingo2audiolanguage, Audioset, chen_vggsound_2020, Librispeech, yang2024airbenchbenchmarkinglargeaudiolanguage] address speech-related tasks (ASR, emotion recognition, and entity recognition) and non-speech tasks (such as sound classification and audio grounding). However, while attempts at creating audio-visual datasets date back to at least 2021 [yun_pano-avqa_2021], these early efforts often had significant limitations. For example, Music-AVQA [li2022learninganswerquestionsdynamic] focused specifically on music performance videos, while Pano-AVQA [yun_pano-avqa_2021] centered on panoramic videos. Others, such as AVQA [avqa] and OmniBench [li_omnibench_2024], were restricted to short, simple videos or static images. Furthermore, AV-Odyssey [gong_av-odyssey_2024] heavily emphasized specific audio tasks, such as recognizing timbre and loudness, rather than broader audio-visual understanding. While WorldSense [hong_worldsense_2025], a concurrent work, also provides a valuable benchmark for real-world audio-visual question-answering, the efficient scalability of AVQA datasets and the enhancement of OLM abilities still require further exploration.
@@ 3 Daily-Omni
This section describes the construction of Daily-Omni, including data curation, video annotation, and QA synthesis and verification. We also describe an evaluation protocol and include a training-free modular reference baseline to help interpret model behaviors. Daily-Omni contains 684 real-world videos from all 11 YouTube categories and 1,197 multiple-choice questions spanning 6 task families. The dataset includes 550 questions on 60-second clips and 647 questions on 30-second clips. Table 1 compares Daily-Omni with prior audio-visual benchmarks.
(a) Word cloud of Daily-Omni QA pairs
(b) Statistics and distributions of the benchmark
Figure 2: Overview of the Daily-Omni benchmark. (a) The prominent vocabulary in the word cloud highlights our focus on rich human activities, diverse acoustic events, and temporal reasoning. (b) The bar chart details the long-tail distribution of the 14 source video categories (top) and the composition of the 6 core question types (bottom).
@@ 3.1 Data Curation
We curate videos with rich temporal dynamics in vision and diverse everyday sounds beyond speech (e.g., music and environmental events). We exclude largely static talking-head videos and non-English videos to reduce confounds from multilingual content. Figure 2 shows the word cloud of questions and choices in Daily-Omni and the distribution of video categories and QA types. Our videos are sourced from AudioSet [Audioset], Video-MME [fu2024videommefirstevercomprehensiveevaluation], and FineVideo [Farre2024FineVideo]. For AudioSet, we retrieve the original YouTube videos and extract 30s/60s segments that contain the target sound events. We use Whisper-V3-Large [radford2022robustspeechrecognitionlargescale] to ensure spoken content is present. For Video-MME and FineVideo, we prioritize clips with substantial audio information and salient temporal visual changes to increase scenario diversity.